# Deterministic Dynamic Maximal Matching in Sublinear Update Time

### v1

@@ 1 Introduction
    the maximum degree in $G[V\setminus V^{\star}]$ is $o(n)$ [^3] Technically, we actually show that $G[V\setminus V^{\star}]$ has arboricity $o(n)$., and
@@
Surprisingly, we can obtain $V^{\star}$ from an edge-degree constrained subgraph (EDCS), a well-known subgraph $H\subseteq G$ that is useful for $(1.5+\epsilon)$-approximate maximum matching algorithms in the dynamic setting and beyond [BS15b, BS16b, AB19, ABB+19a, GSSU22b, AB23, Beh23a, BKS23b, Ber24]. See Definition 2.1 for definition. The outline below presents a completely novel application of this central object in the literature.
The set $V^{\star}$ is simply the set of vertices with “high degree” in the EDCS $H$, which can be explicitly maintained in $o(n)$ time using existing results [GSSU22a]. By the structure of EDCS, Item 1 follows directly. To see why Item 2 should hold, we observe that the graph $H_{\texttt{hilo}}:=E_{H}(V^{\star},V\setminus V^{\star})$ has a degree gap. More precisely, for some number $X$ and $\gamma>0$, every vertex in $V^{\star}$ has degree at least $X$ in $H_{\texttt{hilo}}$, while every vertex in $V\setminus V^{\star}$ has degree at most $(1-\gamma)X$ in $H_{\texttt{hilo}}$. Consequently, for every set $S\subseteq V^{\star}$, we have $|N_{H_{\texttt{hilo}}}(S)|\geq(1+\Omega(\gamma))|S|$, i.e., Hall’s condition holds with a slack. This strong expansion implies the existence of short $O(\log(n)/\gamma)$-length augmenting paths and allows us to maintain a $V^{\star}$-perfect matching in $o(n)$ update time.
@@
We first present the preliminaries in Section 2 and the subroutine for efficiently matching most vertices with almost maximum degree in Section 3. They are needed for our detailed overview in Section 4. Then, we present a complete algorithm for maintaining perfect matching of high-degree vertices in a graph with degree gap in Section 5 and use it to formally prove our main result in Section 6.
@@ 4 Overview: Our Dynamic Maximal Matching Algorithm
For ease of exposition, we first present the algorithm in a decremental setting, where the input graph only undergoes edge deletions. We will show how to handle fully dynamic updates without increasing update time at the end in Section 4.7.
@@
The reader should think of $\epsilon=1/n^{\beta}$ and $B=1/\epsilon^{2}$ for some absolute constant $10<\beta<$ 1.
Let $G=(V,E)$ be the input graph with $n$ nodes, undergoing edge deletions. We maintain a $(B,(1-\epsilon)B)$-EDCS $H:=(V,E(H))$ of $G$ at all times, as per Theorem 2.2. This incurs an update time of $O\left(\frac{n}{B\epsilon}\right)$. The EDCS $H$ is maintained explicitly, and we can make adjacency-list and adjacency-matrix queries to $H$ whenever we want. We maintain the adjacency-lists of all vertices in $H$ as binary search trees. Hence, we may answer an adjacency-list and adjacency-matrix query to $H$ in $O(\log n)$ and $O(1)$ time respectively.
@@
In Section 4.1, we present a classification of nodes based on their degrees in the EDCS $H$ at the start of a phase, under one simplifying assumption (see Assumption 4.1). To highlight the main ideas, we first describe and analyze our algorithm under this assumption. Section 4.2 shows how to maintain a base matching $M_{\texttt{base}}$, which matches all the nodes of a certain type (that have large EDCS degrees). Next, Section 4.3 shows how to maintain a maximal matching in the subgraph of $G$ induced by the nodes that are free under $M_{\texttt{base}}$. In Section 4.4, we put together the algorithms from Sections 4.2 and 4.3, and derive Theorem 1.1. Next, in Section 4.5, we explain how to deal with the scenario when the simplifying assumption we made earlier (see Assumption 4.1) does not hold. Section 4.6 sketches the proof outline of a key lemma, which allows us to maintain the matching $M_{\texttt{base}}$ efficiently in Section 4.2. Finally, Section 4.7 shows how to extend our algorithm so that it can deal with fully dynamic updates.
@@ 4.1 Classification of Nodes and a Simplifying Assumption
We will outline how to maintain a maximal matching $M_{\texttt{final}}$ in $G$ during a phase. We start by making one simplifying assumption stated below. Towards the end of this section, we explain how to adapt our algorithm when this assumption does not hold (see Section 4.5).
@@
In other words, this establishes a small “degree-gap” within the EDCS $H$ at the start of the phase, by asserting that there does not exist any node $v\in V$ with $\left(\frac{1}{2}-\delta\right)B\leq\mathrm{deg}_{H_{\texttt{init}}}(v)\leq\left(\frac{1}{2}+\delta\right)B$. We say that a node $v\in V$ is very high if $\mathrm{deg}_{H_{\texttt{init}}}(v)>\left(\frac{1}{2}+\delta\right)B$, and low if $\mathrm{deg}_{H_{\texttt{init}}}(v)<\left(\frac{1}{2}-\delta\right)B$. We let $V_{\texttt{v-hi}}$ and $V_{\texttt{lo}}$ respectively denote the sets of very high and low nodes. By Assumption 4.1, the node-set $V$ is partitioned into $V_{\texttt{v-hi}}\subseteq V$ and $V_{\texttt{lo}}=V\setminus V_{\texttt{v-hi}}$. Our algorithm will crucially use two properties that arise out of this classification of nodes, as described below.
@@
First, we prove Item 1. Note that since $(u,v)\in E(H_{\texttt{init}})$, we have $\mathrm{deg}_{H_{\texttt{init}}}(u)+\mathrm{deg}_{H_{\texttt{init}}}(v)\leq B$ (see Definition 2.1). Furthermore, since $u$ is very-high, we have $\mathrm{deg}_{H_{\texttt{init}}}(u)>\left(\frac{1}{2}+\delta\right)B$. Thus, it follows that $\mathrm{deg}_{H_{\texttt{init}}}(v)\leq B-\mathrm{deg}_{H_{\texttt{init}}}(u)<\left(\frac{1}{2}-\delta\right)B$, and so $v$ must be a low node.
Next, we prove Item 2. Let $e=(u,v)$, where both $u,v$ are low nodes. Then $\mathrm{deg}_{H_{\texttt{init}}}(u)<\left(\frac{1}{2}-\delta\right)B$ and $\mathrm{deg}_{H_{\texttt{init}}}(v)<\left(\frac{1}{2}-\delta\right)B$, and hence $\mathrm{deg}_{H_{\texttt{init}}}(u)+\mathrm{deg}_{H_{\texttt{init}}}(v)<(1-2\delta)B<(1-\epsilon)B$. As $H_{\texttt{init}}$ is a $(B,(1-\epsilon)B)$-EDCS of $G_{\texttt{init}}$, it follows that $(u,v)\in E\left(H_{\texttt{init}}\right)$ (see Definition 2.1). ∎
@@ 4.2 Maintaining a Matching of All the Safe Nodes
One of our key technical insights is that the degree gap in $H_{\texttt{core}}$ allows us to efficiently maintain a matching $M_{\texttt{base}}\subseteq E(H_{\texttt{core}})$ that matches all safe nodes, as per Lemma 4.5 below. With appropriate setting of parameters, the update time of both of the algorithms (deterministic and randomized) in Lemma 4.5 are sublinear in $n$.
@@
The only property of $H_{\texttt{core}}$ that we need for Lemma Lemma 4.5 is the degree gap between $V_{\texttt{sf}}$ and $V_{\texttt{lo}}$; in fact, the whole reason we use an EDCS in the first place is to ensure this degree gap.
We defer the proof sketch of Lemma 4.5 to Section 4.6, but to provide an intuition for why the degree gap helps match vertices in $V_{\texttt{sf}}$, we show here that it is easy to deterministically maintain a fractional matching satisfying the concerned properties of Lemma 4.5, in $O(B/\delta)$ update time. We do not this result in the main body of our paper and include it only for intuition. We denote the fractional matching by $x_{\texttt{base}}:E(H_{\texttt{core}})\rightarrow[0,1]$, which is constructed as follows.
{mdframed}
[hidealllines=true, backgroundcolor=gray!20, leftmargin=0cm,innerleftmargin=0.4cm,innerrightmargin=0.4cm,innertopmargin=0.4cm,innerbottommargin=0.4cm,roundcorner=10pt]
  • $\Gamma\leftarrow\left(\frac{1}{2}+\delta-2\epsilon\right)B$ and $x_{\texttt{base}}(e)\leftarrow 0$ for all $e\in E$
  • For every safe node $v\in V_{\texttt{sf}}$
    • Let $E^{\star}_{v}\subseteq E(H_{\texttt{core}})$ be a set of $\Gamma$ distinct edges incident to $v$. ($E^{\star}_{v}$ exists as $v\in V_{\texttt{sf}}$).
    • $x_{\texttt{base}}(e)\leftarrow 1/\Gamma$ for all $e\in E^{\star}_{v}$
  • Return $x_{\texttt{base}}:E(H_{\texttt{core}})\rightarrow[0,1]$.
Consider any two distinct nodes $v,v^{\prime}\in V_{\texttt{sf}}\subseteq V_{\texttt{v-hi}}$. By Item 1 of Proposition 4.2, there cannot be an edge $(v,v^{\prime})\in E(H_{\texttt{core}})\subseteq E(H_{\texttt{init}})$, and hence $E^{\star}_{v}\cap E^{\star}_{v^{\prime}}=\emptyset$. Using this observation, it is easy to verify that the function $x_{\texttt{base}}$ is a valid fractional matching in $H_{\texttt{core}}$, and that every node $v\in V_{\texttt{sf}}$ receives a weight $x_{\texttt{base}}(v)=1$ under this fractional matching. [^4] The weight $x_{\texttt{base}}(v)$ is the sum of the values $x_{\texttt{base}}(e)$, over all the edges in $H_{\texttt{core}}$ that are incident on $v$. Furthermore, since the EDCS $H_{\texttt{init}}$ has maximum degree at most $B$ (this follows from Definition 2.1) and each phase lasts for $\delta n$ updates, we can deterministically maintain the fractional matching $x_{\texttt{base}}$ in $O\left(\frac{nB}{\delta n}\right)=O(B/\delta)$ amortized update time. One can also verify that each update in $G$ that is internal to a phase leads to at most $O(1)$ nodes $v\in V$ changing their weights $x_{\texttt{base}}(v)$ under $x_{\texttt{base}}$.
@@ 4.3 Maintaining a Maximal Matching in the Adjunct Subgraph
Suppose that we maintain the matching $M_{\texttt{base}}$ as per Lemma 4.5. Then with an additive overhead of $\tilde{O}\left(\frac{B}{\delta}+\frac{n}{\delta B}\right)$ amortized update time, we can deterministically and explicitly maintain a maximal matching $M_{\texttt{adj}}\subseteq E_{\texttt{adj}}$ in $G_{\texttt{adj}}$.
At this point, the reader might get alarmed because $G_{\texttt{adj}}$ undergoes node-updates every time an edge gets added to/removed from $M_{\texttt{base}}$, which might indicate that it is impossible to efficiently maintain a maximal matching in $G_{\texttt{adj}}$. To assuage this concern, we emphasize that: (i) the arboricity of $G_{\texttt{adj}}$ is sublinear in $n$ (see Claim 4.8), and (ii) an edge update in $G$ that is internal to a phase leads to at most $O(1)$ node-updates in $G_{\texttt{adj}}$ (see Claim 4.9). These two properties ensure that we can maintain the matching $M_{\texttt{adj}}$ in sublinear in $n$ update time.
@@
Since $V_{\texttt{adj}}:=V\setminus V(M_{\texttt{base}})$Item 1 follows from Item 1 of Lemma 4.5, and Corollary 4.4.
To prove Item 2, consider any node $v\in V_{\texttt{lo}}$. Let $u\in V_{\texttt{lo}}$ be any low neighbor of $v$ in $E_{\texttt{adj}}\subseteq E\subseteq E_{\texttt{init}}$. Then, by Item 2 of Proposition 4.2, we have $(u,v)\in H_{\texttt{init}}$. Accordingly, the number of low neighbors of $v$ in $G_{\texttt{adj}}$ is at most $\mathrm{deg}_{H_{\texttt{init}}}(v)\leq B$, where the last inequality holds because $H_{\texttt{init}}$ is a $(B,(1-\epsilon)B)$-EDCS of $G_{\texttt{init}}$ (see Definition 2.1). As the node-set $V$ is partitioned into $V_{\texttt{lo}},V_{\texttt{v-hi}}$ (see Assumption 4.1) and $V_{\texttt{v-hi}}$ is further partitioned into $V_{\texttt{sf}},V_{\texttt{dmg}}$, we get:
@@
where the last equality follows from Item 1. ∎
@@
Since $V_{\texttt{adj}}:=V\setminus V(M_{\texttt{base}})$, the claim follows from Item 2 of Lemma 4.5. ∎
@@
By Proposition 4.2(2) we have that $E_{\texttt{adj}}\subseteq H_{\texttt{init}}$. As $H_{\texttt{init}}$ is a $(B,B(1-\epsilon))$-EDCS and the maximum vertex degree of a $(B,B(1-\epsilon))$-EDCS is $B$ we must have that $|E_{\texttt{adj}}|\leq|H_{\texttt{init}}|=O(nB)$.
Proof of Lemma 4.7
At the start of a phase, we initialize $M_{\texttt{adj}}$ to be any arbitrary maximal matching in $G_{\texttt{adj}}$. By Corollary 4.10, this takes $O(Bn)$ time. Since the phase lasts for $\delta n$ updates in $G$, this step incurs an amortized update time of:
@@
Subsequently, to handle the updates during a phase, we use the following auxiliary data structures: Each node $v\in V_{\texttt{dmg}}$ maintains the set $\mathcal{F}_{\texttt{adj}}(v):=\left\{u\in V_{\texttt{adj}}\setminus V(M_{\texttt{adj}}):(u,v)\in E_{\texttt{adj}}\right\}$ of its free neighbors (in $G_{\texttt{adj}}$) under the matching $M_{\texttt{adj}}$. [^5] Maximality of $M_{\texttt{adj}}$ implies that $\mathcal{F}_{\texttt{adj}}(v)=\emptyset$ for all $v\in V_{\texttt{dmg}}\setminus V(M_{\texttt{adj}})$. Whenever a node $v$ moves from $V_{\texttt{sf}}$ to $V_{\texttt{dmg}}$, we spend $\tilde{O}(n)$ time to initialize the set $\mathcal{F}_{\texttt{adj}}(v)$ as a balanced search tree. Within a phase, a node can move from $V_{\texttt{sf}}$ to $V_{\texttt{dmg}}$ at most once (see Definition 4.3 and Corollary 4.4). Thus, by Item 1 of Claim 4.8, these initializations take $\tilde{O}\left(\frac{n}{B}\cdot n\right)=\tilde{O}\left(\frac{n^{2}}{B}\right)$ total time during a phase. As each phase lasts for $\delta n$ updates in $G$, this incurs an amortized update time of $\tilde{O}\left(\frac{n^{2}}{B\cdot\delta n}\right)=\tilde{O}\left(\frac{n}{\delta B}\right)$.
Claim 4.9 guarantees that each edge update in $G$ leads to $O(1)$ node-updates in $G_{\texttt{adj}}$. We now show that to maintain the maximal matching $M_{\texttt{adj}}$, a node-update to $G_{\texttt{adj}}$ can be handled in $\tilde{O}\left(B+\frac{n}{B}\right)$ worst-case time. If a new vertex $v$ enters $G_{\texttt{adj}}$ (because $v$ became unmatched in $M_{\texttt{base}}$), then to maintain the maximality of $M_{\texttt{adj}}$, the algorithm needs to find a free neighbor of $v$ (if one exists). The same needs to be done if $x$ leaves $G_{\texttt{adj}}$, and $x$ was previously matched to $v$ in $M_{\texttt{adj}}$. So to handle a node-update in $G_{\texttt{adj}}$ we need a subroutine that takes in a vertex $v$ and either finds a $M_{\texttt{adj}}$-free neighbor of $v$ in $G_{\texttt{adj}}$ or certifies that none exists.
The above subroutine can be performed in $\tilde{O}(1)$ time if $v\in V_{\texttt{dmg}}$ (using the set $\mathcal{F}_{\texttt{adj}}(v)$); if $v\in V_{\texttt{adj}}\setminus V_{\texttt{dmg}}$ then by Item 1 of Claim 4.8 $v\in V_{\texttt{lo}}$, so by Item 2 of the same claim we can perform the subroutine in time $\mathrm{deg}_{G_{\texttt{adj}}}(v)=O\left(B+\frac{n}{B}\right)$. Further, once we decide to match a node $u\in V_{\texttt{adj}}$ to one of its free neighbors, we can update all the relevant sets $\mathcal{F}_{\texttt{adj}}(v)$ in $\tilde{O}\left(\left|V_{\texttt{dmg}}\right|\right)=\tilde{O}\left(\frac{n}{B}\right)$ time (see Claim 4.8).

### v2

@@ 1 Introduction
    the maximum degree in $G[V\setminus V^{\star}]$ is o(n)o(n) [^3] Technically, we actually show that $G[V\setminus V^{\star}]$ has arboricity $o(n)$., and
@@
Surprisingly, we can obtain $V^{\star}$ from an edge-degree constrained subgraph (EDCS), a well-known subgraph $H\subseteq G$ that is useful for $(1.5+\epsilon)$-approximate maximum matching algorithms in the dynamic setting and beyond [BS15b, BS16b, AB19, ABB+19a, GSSU22b, AB23, Beh23a, BKS23b, Ber24]. See Definition˜2.1 for definition. The outline below presents a completely novel application of this central object in the literature.
The set $V^{\star}$ is simply the set of vertices with “high degree” in the EDCS $H$, which can be explicitly maintained in $o(n)$ time using existing results [GSSU22a]. By the structure of EDCS, Item˜1 follows directly. To see why Item˜2 should hold, we observe that the graph $H_{\texttt{hilo}}:=E_{H}(V^{\star},V\setminus V^{\star})$ has a degree gap. More precisely, for some number $X$ and $\gamma>0$, every vertex in $V^{\star}$ has degree at least $X$ in $H_{\texttt{hilo}}$, while every vertex in $V\setminus V^{\star}$ has degree at most $(1-\gamma)X$ in $H_{\texttt{hilo}}$. Consequently, for every set $S\subseteq V^{\star}$, we have $|N_{H_{\texttt{hilo}}}(S)|\geq(1+\Omega(\gamma))|S|$, i.e., Hall’s condition holds with a slack. This strong expansion implies the existence of short $O(\log(n)/\gamma)$-length augmenting paths and allows us to maintain a $V^{\star}$-perfect matching in $o(n)$ update time.
@@
We first present the preliminaries in Section˜2 and the subroutine for efficiently matching most vertices with almost maximum degree in Section˜3. They are needed for our detailed overview in Section˜4. Then, we present a complete algorithm for maintaining perfect matching of high-degree vertices in a graph with degree gap in Section˜5 and use it to formally prove our main result in Section˜6.
@@ 4 Overview: Our Dynamic Maximal Matching Algorithm
For ease of exposition, we first present the algorithm in a decremental setting, where the input graph only undergoes edge deletions. We will show how to handle fully dynamic updates without increasing update time at the end in Section˜4.7.
@@
The reader should think of $\epsilon=1/n^{\beta}$ and $B=1/\epsilon^{2}$ for some absolute constant $0<\beta<1$.
Let $G=(V,E)$ be the input graph with $n$ nodes, undergoing edge deletions. We maintain a $(B,(1-\epsilon)B)$-EDCS $H:=(V,E(H))$ of $G$ at all times, as per Theorem˜2.2. This incurs an update time of $O\left(\frac{n}{B\epsilon}\right)$. The EDCS $H$ is maintained explicitly, and we can make adjacency-list and adjacency-matrix queries to $H$ whenever we want. We maintain the adjacency-lists of all vertices in $H$ as binary search trees. Hence, we may answer an adjacency-list and adjacency-matrix query to $H$ in $O(\log n)$ and $O(1)$ time respectively.
@@
In Section˜4.1, we present a classification of nodes based on their degrees in the EDCS $H$ at the start of a phase, under one simplifying assumption (see Assumption˜4.1). To highlight the main ideas, we first describe and analyze our algorithm under this assumption. Section˜4.2 shows how to maintain a base matching $M_{\texttt{base}}$, which matches all the nodes of a certain type (that have large EDCS degrees). Next, Section˜4.3 shows how to maintain a maximal matching in the subgraph of $G$ induced by the nodes that are free under $M_{\texttt{base}}$. In Section˜4.4, we put together the algorithms from Sections˜4.2 and 4.3, and derive Theorem˜1.1. Next, in Section˜4.5, we explain how to deal with the scenario when the simplifying assumption we made earlier (see Assumption˜4.1) does not hold. Section˜4.6 sketches the proof outline of a key lemma, which allows us to maintain the matching $M_{\texttt{base}}$ efficiently in Section˜4.2. Finally, Section˜4.7 shows how to extend our algorithm so that it can deal with fully dynamic updates.
@@ 4.1 Classification of Nodes and a Simplifying Assumption
We will outline how to maintain a maximal matching $M_{\texttt{final}}$ in $G$ during a phase. We start by making one simplifying assumption stated below. Towards the end of this section, we explain how to adapt our algorithm when this assumption does not hold (see Section˜4.5).
@@
In other words, this establishes a small “degree-gap” within the EDCS $H$ at the start of the phase, by asserting that there does not exist any node $v\in V$ with $\left(\frac{1}{2}-\delta\right)B\leq\mathrm{deg}_{H_{\texttt{init}}}(v)\leq\left(\frac{1}{2}+\delta\right)B$. We say that a node $v\in V$ is very high if $\mathrm{deg}_{H_{\texttt{init}}}(v)>\left(\frac{1}{2}+\delta\right)B$, and low if $\mathrm{deg}_{H_{\texttt{init}}}(v)<\left(\frac{1}{2}-\delta\right)B$. We let $V_{\texttt{v-hi}}$ and $V_{\texttt{lo}}$ respectively denote the sets of very high and low nodes. By Assumption˜4.1, the node-set $V$ is partitioned into $V_{\texttt{v-hi}}\subseteq V$ and $V_{\texttt{lo}}=V\setminus V_{\texttt{v-hi}}$. Our algorithm will crucially use two properties that arise out of this classification of nodes, as described below.
@@
First, we prove Item˜1. Note that since $(u,v)\in E(H_{\texttt{init}})$, we have $\mathrm{deg}_{H_{\texttt{init}}}(u)+\mathrm{deg}_{H_{\texttt{init}}}(v)\leq B$ (see Definition˜2.1). Furthermore, since $u$ is very-high, we have $\mathrm{deg}_{H_{\texttt{init}}}(u)>\left(\frac{1}{2}+\delta\right)B$. Thus, it follows that $\mathrm{deg}_{H_{\texttt{init}}}(v)\leq B-\mathrm{deg}_{H_{\texttt{init}}}(u)<\left(\frac{1}{2}-\delta\right)B$, and so $v$ must be a low node.
Next, we prove Item˜2. Let $e=(u,v)$, where both $u,v$ are low nodes. Then $\mathrm{deg}_{H_{\texttt{init}}}(u)<\left(\frac{1}{2}-\delta\right)B$ and $\mathrm{deg}_{H_{\texttt{init}}}(v)<\left(\frac{1}{2}-\delta\right)B$, and hence $\mathrm{deg}_{H_{\texttt{init}}}(u)+\mathrm{deg}_{H_{\texttt{init}}}(v)<(1-2\delta)B<(1-\epsilon)B$. As $H_{\texttt{init}}$ is a $(B,(1-\epsilon)B)$-EDCS of $G_{\texttt{init}}$, it follows that $(u,v)\in E\left(H_{\texttt{init}}\right)$ (see Definition˜2.1). ∎
@@ 4.2 Maintaining a Matching of All the Safe Nodes
One of our key technical insights is that the degree gap in $H_{\texttt{core}}$ allows us to efficiently maintain a matching $M_{\texttt{base}}\subseteq E(H_{\texttt{core}})$ that matches all safe nodes, as per Lemma˜4.5 below. With appropriate setting of parameters, the update time of both of the algorithms (deterministic and randomized) in Lemma˜4.5 are sublinear in $n$.
@@
The only property of $H_{\texttt{core}}$ that we need for Lemma Lemma˜4.5 is the degree gap between $V_{\texttt{sf}}$ and $V_{\texttt{lo}}$; in fact, the whole reason we use an EDCS in the first place is to ensure this degree gap.
We defer the proof sketch of Lemma˜4.5 to Section˜4.6, but to provide an intuition for why the degree gap helps match vertices in $V_{\texttt{sf}}$, we show here that it is easy to deterministically maintain a fractional matching satisfying the concerned properties of Lemma˜4.5, in $O(B/\delta)$ update time. We do not this result in the main body of our paper and include it only for intuition. We denote the fractional matching by $x_{\texttt{base}}:E(H_{\texttt{core}})\rightarrow[0,1]$, which is constructed as follows.
$\Gamma\leftarrow\left(\frac{1}{2}+\delta-2\epsilon\right)B$ and $x_{\texttt{base}}(e)\leftarrow 0$ for all $e\in E$ For every safe node $v\in V_{\texttt{sf}}$ Let $E^{\star}_{v}\subseteq E(H_{\texttt{core}})$ be a set of $\Gamma$ distinct edges incident to $v$. ($E^{\star}_{v}$ exists as $v\in V_{\texttt{sf}}$). $x_{\texttt{base}}(e)\leftarrow 1/\Gamma$ for all $e\in E^{\star}_{v}$ Return $x_{\texttt{base}}:E(H_{\texttt{core}})\rightarrow[0,1]$.
Consider any two distinct nodes $v,v^{\prime}\in V_{\texttt{sf}}\subseteq V_{\texttt{v-hi}}$. By Item˜1 of Proposition˜4.2, there cannot be an edge $(v,v^{\prime})\in E(H_{\texttt{core}})\subseteq E(H_{\texttt{init}})$, and hence $E^{\star}_{v}\cap E^{\star}_{v^{\prime}}=\emptyset$. Using this observation, it is easy to verify that the function $x_{\texttt{base}}$ is a valid fractional matching in $H_{\texttt{core}}$, and that every node $v\in V_{\texttt{sf}}$ receives a weight $x_{\texttt{base}}(v)=1$ under this fractional matching. [^4] The weight $x_{\texttt{base}}(v)$ is the sum of the values $x_{\texttt{base}}(e)$, over all the edges in $H_{\texttt{core}}$ that are incident on $v$. Furthermore, since the EDCS $H_{\texttt{init}}$ has maximum degree at most $B$ (this follows from Definition˜2.1) and each phase lasts for $\delta n$ updates, we can deterministically maintain the fractional matching $x_{\texttt{base}}$ in $O\left(\frac{nB}{\delta n}\right)=O(B/\delta)$ amortized update time. One can also verify that each update in $G$ that is internal to a phase leads to at most $O(1)$ nodes $v\in V$ changing their weights $x_{\texttt{base}}(v)$ under $x_{\texttt{base}}$.
@@ 4.3 Maintaining a Maximal Matching in the Adjunct Subgraph
Suppose that we maintain the matching $M_{\texttt{base}}$ as per Lemma˜4.5. Then with an additive overhead of $\tilde{O}\left(\frac{B}{\delta}+\frac{n}{\delta B}\right)$ amortized update time, we can deterministically and explicitly maintain a maximal matching $M_{\texttt{adj}}\subseteq E_{\texttt{adj}}$ in $G_{\texttt{adj}}$.
At this point, the reader might get alarmed because $G_{\texttt{adj}}$ undergoes node-updates every time an edge gets added to/removed from $M_{\texttt{base}}$, which might indicate that it is impossible to efficiently maintain a maximal matching in $G_{\texttt{adj}}$. To assuage this concern, we emphasize that: (i) the arboricity of $G_{\texttt{adj}}$ is sublinear in $n$ (see Claim˜4.8), and (ii) an edge update in $G$ that is internal to a phase leads to at most $O(1)$ node-updates in $G_{\texttt{adj}}$ (see Claim˜4.9). These two properties ensure that we can maintain the matching $M_{\texttt{adj}}$ in sublinear in $n$ update time.
@@
Since $V_{\texttt{adj}}:=V\setminus V(M_{\texttt{base}})$Item˜1 follows from Item˜1 of Lemma˜4.5, and Corollary˜4.4.
To prove Item˜2, consider any node $v\in V_{\texttt{lo}}$. Let $u\in V_{\texttt{lo}}$ be any low neighbor of $v$ in $E_{\texttt{adj}}\subseteq E\subseteq E_{\texttt{init}}$. Then, by Item˜2 of Proposition˜4.2, we have $(u,v)\in H_{\texttt{init}}$. Accordingly, the number of low neighbors of $v$ in $G_{\texttt{adj}}$ is at most $\mathrm{deg}_{H_{\texttt{init}}}(v)\leq B$, where the last inequality holds because $H_{\texttt{init}}$ is a $(B,(1-\epsilon)B)$-EDCS of $G_{\texttt{init}}$ (see Definition˜2.1). As the node-set $V$ is partitioned into $V_{\texttt{lo}},V_{\texttt{v-hi}}$ (see Assumption˜4.1) and $V_{\texttt{v-hi}}$ is further partitioned into $V_{\texttt{sf}},V_{\texttt{dmg}}$, we get:
@@
where the last equality follows from Item˜1. ∎
@@
Since $V_{\texttt{adj}}:=V\setminus V(M_{\texttt{base}})$, the claim follows from Item˜2 of Lemma˜4.5. ∎
@@
By Proposition˜4.2(2) we have that $E_{\texttt{adj}}\subseteq H_{\texttt{init}}$. As $H_{\texttt{init}}$ is a $(B,B(1-\epsilon))$-EDCS and the maximum vertex degree of a $(B,B(1-\epsilon))$-EDCS is $B$ we must have that $|E_{\texttt{adj}}|\leq|H_{\texttt{init}}|=O(nB)$.
Proof of Lemma˜4.7
At the start of a phase, we initialize $M_{\texttt{adj}}$ to be any arbitrary maximal matching in $G_{\texttt{adj}}$. By Corollary˜4.10, this takes $O(Bn)$ time. Since the phase lasts for $\delta n$ updates in $G$, this step incurs an amortized update time of:
@@
Subsequently, to handle the updates during a phase, we use the following auxiliary data structures: Each node $v\in V_{\texttt{dmg}}$ maintains the set $\mathcal{F}_{\texttt{adj}}(v):=\left\{u\in V_{\texttt{adj}}\setminus V(M_{\texttt{adj}}):(u,v)\in E_{\texttt{adj}}\right\}$ of its free neighbors (in $G_{\texttt{adj}}$) under the matching $M_{\texttt{adj}}$. [^5] Maximality of $M_{\texttt{adj}}$ implies that $\mathcal{F}_{\texttt{adj}}(v)=\emptyset$ for all $v\in V_{\texttt{dmg}}\setminus V(M_{\texttt{adj}})$. Whenever a node $v$ moves from $V_{\texttt{sf}}$ to $V_{\texttt{dmg}}$, we spend $\tilde{O}(n)$ time to initialize the set $\mathcal{F}_{\texttt{adj}}(v)$ as a balanced search tree. Within a phase, a node can move from $V_{\texttt{sf}}$ to $V_{\texttt{dmg}}$ at most once (see Definition˜4.3 and Corollary˜4.4). Thus, by Item˜1 of Claim˜4.8, these initializations take $\tilde{O}\left(\frac{n}{B}\cdot n\right)=\tilde{O}\left(\frac{n^{2}}{B}\right)$ total time during a phase. As each phase lasts for $\delta n$ updates in $G$, this incurs an amortized update time of $\tilde{O}\left(\frac{n^{2}}{B\cdot\delta n}\right)=\tilde{O}\left(\frac{n}{\delta B}\right)$.
Claim˜4.9 guarantees that each edge update in $G$ leads to $O(1)$ node-updates in $G_{\texttt{adj}}$. We now show that to maintain the maximal matching $M_{\texttt{adj}}$, a node-update to $G_{\texttt{adj}}$ can be handled in $\tilde{O}\left(B+\frac{n}{B}\right)$ worst-case time. If a new vertex $v$ enters $G_{\texttt{adj}}$ (because $v$ became unmatched in $M_{\texttt{base}}$), then to maintain the maximality of $M_{\texttt{adj}}$, the algorithm needs to find a free neighbor of $v$ (if one exists). The same needs to be done if $x$ leaves $G_{\texttt{adj}}$, and $x$ was previously matched to $v$ in $M_{\texttt{adj}}$. So to handle a node-update in $G_{\texttt{adj}}$ we need a subroutine that takes in a vertex $v$ and either finds a $M_{\texttt{adj}}$-free neighbor of $v$ in $G_{\texttt{adj}}$ or certifies that none exists.
The above subroutine can be performed in $\tilde{O}(1)$ time if $v\in V_{\texttt{dmg}}$ (using the set $\mathcal{F}_{\texttt{adj}}(v)$); if $v\in V_{\texttt{adj}}\setminus V_{\texttt{dmg}}$ then by Item˜1 of Claim˜4.8 $v\in V_{\texttt{lo}}$, so by Item˜2 of the same claim we can perform the subroutine in time $\mathrm{deg}_{G_{\texttt{adj}}}(v)=O\left(B+\frac{n}{B}\right)$. Further, once we decide to match a node $u\in V_{\texttt{adj}}$ to one of its free neighbors, we can update all the relevant sets $\mathcal{F}_{\texttt{adj}}(v)$ in $\tilde{O}\left(\left|V_{\texttt{dmg}}\right|\right)=\tilde{O}\left(\frac{n}{B}\right)$ time (see Claim˜4.8).